# SDDBench: A Benchmark for Synthesizable Drug Design

### v1

@@ 6 Conclusion
In this work, we propose a novel round-trip score to assess the synthesizability of molecules generated by existing SBDD models. This score evaluates the likelihood that a retrosynthetic planner, trained on current reaction data, can predict feasible synthetic routes for these molecules. To enhance the robustness of our evaluation method, we advocate for the release of additional reaction data. This expanded dataset would significantly improve the accuracy and reliability of our assessments.
Appendix A Dataset
A.1 Intended Use
This paper introduces a dataset and benchmark designed to evaluate the round-trip score of molecules predicted by drug design generative models. The round-trip score assesses the likelihood of finding a feasible synthetic route for a predicted molecule using retrosynthetic planners trained on current reaction data. To calculate this score, we train both a retrosynthetic planner and a reaction prediction model using the open-source USPTO-full reaction dataset.
Our provided code repository includes pre-trained model files for both the retrosynthesis prediction model (Neuralsym) and the reaction prediction model (based on a Transformer Decoder-only architecture). Neuralsym, combined with beam search, serves as our retrosynthetic planner.
We evaluate the round-trip scores for molecules generated by various drug design models across 100 protein targets from the Cross-Docked test set. The generated molecule files are obtained from the official GitHub repositories associated with each generative model’s publication.
A.2 Dataset Details
We begin by cleaning the USPTO-full dataset and removing duplicate reactions. The dataset is then split, with 90% used for training and 10% for testing the reaction prediction model. For the retrosynthesis prediction model, we take a different approach. Instead of using all processed reaction data, we construct a reaction network from the cleaned data and extract synthetic routes, ensuring that the leaf nodes of valid routes are starting materials. These synthetic routes are then split based on molecules: 98% for training, 1% for validation, and 1% for testing. Detailed information about the data used to train both the retrosynthesis prediction model and the reaction prediction model can be found in Tables 2 and 3.
| 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 | |
| Training | 56,626 | 24,933 | 12,566 | 5,853 | 3,168 | 1,259 | 512 | 191 | 85 | 17 | 7 | 0 | 0 | 1 |
| Validation | 577 | 254 | 128 | 59 | 32 | 12 | 5 | 1 | 0 | 0 | 0 | 0 | 0 | 0 |
| Test | 577 | 254 | 128 | 59 | 32 | 12 | 5 | 1 | 0 | 0 | 0 | 0 | 0 | 0 |
| Training | 811,602 |
| Test | 90,718 |
A.3 Reproducibility
We use Pytorch (Paszke et al., 2019) to implement our retrosynthesis and reaction prediction models. The softwares that we use for experiments are Python 3.6.8, CUDA 10.2.89, CUDNN 7.6.5, einops 0.4.1, pytorch 1.9.0, pytorch-scatter 2.0.9, pytorch-sparse 0.6.12, numpy 1.19.2, torchvision 0.10.0, and torchdrug 0.1.3.
| max length | 402 |
| embedding size | 64 |
| decoder layers | 6 |
| attention heads | 8 |
| FFN hidden | 2048 |
| dropout | 0.1 |
| epochs | 2000 |
| batch size | 128 |
| warmup | 16000 |
| lr factor | 20 |
| scheduling |
@@ Appendix B Additional Experimental Results
B.1 Synthesizability Evaluation
Tables 5 and 7 present performance gaps between consecutive models for top-k metrics with thresholds of 0.9 and 0.8, respectively. Table 6 compares model performance using top-k route quality with a threshold of 0.8.
@@ B.1 Synthesizability Evaluation
| Pocket2Mol to FLAG | 3.76 | 4.90 | 5.56 | 5.90 | 6.22 |
| FLAG to AR | 2.46 | 3.33 | 3.72 | 4.10 | 4.36 |
| AR to DrugGPS | 0.89 | 0.92 | 0.95 | 0.86 | 0.84 |
| DrugGPS to DecompDiff | 1.58 | 2.10 | 2.42 | 2.53 | 2.76 |
| DecompDiff to TargetDiff | 0.91 | 1.05 | 1.14 | 1.21 | 1.18 |
| TargetDiff to LiGAN | 0.45 | 0.61 | 0.66 | 0.73 | 0.72 |
| Model | Top-1 Max | Top-2 Max | Top-3 Max | Top-4 Max | Top-5 Max |
| CVAE | 1.35 | 1.71 | 1.93 | 2.05 | 2.13 |
| TargetDiff | 1.81 | 2.32 | 2.59 | 2.77 | 2.87 |
| DecompDiff | 2.95 | 3.85 | 4.33 | 4.58 | 4.69 |
| DrugGPS | 4.53 | 5.73 | 6.35 | 6.75 | 7.06 |
| AR | 5.43 | 6.61 | 7.34 | 7.63 | 7.94 |
| FLAG | 8.01 | 10.15 | 11.29 | 11.94 | 12.48 |
| Pocket2Mol | 12.06 | 15.45 | 17.16 | 18.26 | 19.07 |
| Gap Between Models | Top-1 Max | Top-2 Max | Top-3 Max | Top-4 Max | Top-5 Max |
| Pocket2Mol to FLAG | 4.05 | 5.30 | 5.87 | 6.32 | 6.59 |
| FLAG to AR | 2.58 | 3.54 | 3.95 | 4.31 | 4.54 |
| AR to DrugGPS | 0.90 | 0.88 | 0.99 | 0.88 | 0.88 |
| DrugGPS to DecompDiff | 1.58 | 1.88 | 2.02 | 2.17 | 2.37 |
| DecompDiff to TargetDiff | 1.14 | 1.53 | 1.74 | 1.81 | 1.82 |
| TargetDiff to CVAE | 0.46 | 0.61 | 0.66 | 0.72 | 0.74 |

### v2

@@ 6 Conclusion
In this work, we introduce a novel round-trip score to evaluate the molecule synthesizability. Extensive experiments demonstrate that our proposed metric outperforms the search success rate in assessing the feasibility of synthetic routes. We also evaluate the synthesizability of molecules generated by current SBDD models. Future works can adopt our proposed metric for designing synthesizable drugs.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Reproducibility
| Max Length | Embedding Size | Decoder Layers | Attention Heads | FFN Hidden | Dropout |
| 402 | 64 | 6 | 8 | 2048 | 0.1 |
| Epochs | Batch Size | Warmup | LR Factor | Scheduling | |
| 2000 | 128 | 16000 | 20 |
@@ Appendix B Additional Experimental Results
Table 4 presents performance gaps between consecutive models.
@@
| LiGAN to TargetDiff | 0.35 | 0.53 | 0.57 | 0.56 | 0.54 |
| TargetDiff to DecompDiff | 0.03 | 0.33 | 0.48 | 0.66 | 0.64 |
| DecompDiff to DrugGPS | 4.65 | 4.44 | 4.45 | 4.34 | 4.36 |
| DrugGPS to AR | 0.06 | 0.42 | 0.41 | 0.52 | 0.62 |
| AR to FLAG | 5.85 | 6.04 | 6.09 | 6.14 | 6.19 |
| FLAG to Pocket2Mol | 6.05 | 6.41 | 6.72 | 6.83 | 6.83 |